OpenStudy (anonymous):

after 30 hours a sample of Plutonium-243, has decayed to (1/64) of its original mass. find its half life.

OpenStudy (anonymous):

960 hr?

OpenStudy (lgbasallote):

since it;s decay, we use \[\ln (\frac{x_0}{x}) = kt\] after 30 hours x = 1/64 x_0 \[\ln (\frac{x_0}{\frac{1}{64}x_0}) = 30k\] \[\ln (\frac{1}{\frac{1}{64}}) = 30k\] \[\ln (64) = 30k\] \[k = \frac{\ln (64)}{30}\] now to find half-life use \[t_{1/2} = \frac{\ln 2}{k}\] \[t_{1/2} = \frac{\ln 2}{\frac{\ln(64)}{30}}\] \[t_{1/2} = \frac{30\ln 2}{\ln (64)}\] use a calculator to solve this

OpenStudy (anonymous):

thanks